OpenStudy (anonymous):

2)f(x)=4^xlog(small 7)(x)

OpenStudy (ash2326):

@Dodo1 What do you need to do here?

OpenStudy (anonymous):

Logarithmic Derivatives

OpenStudy (anonymous):

f(x)'=

OpenStudy (anonymous):

@Luis_Rivera

jimthompson5910 (jim_thompson5910):

y = 4^(x*log(7,x)) ln(y) = ln( 4^(x*log(7,x)) ) ln(y) = (x*log(7,x))* ln( 4 ) does that help?

OpenStudy (anonymous):

log 7 *x?

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

about part of log (7*x)

OpenStudy (anonymous):

is it times?

jimthompson5910 (jim_thompson5910):

oh log(7,x) means "log base 7 of x"

jimthompson5910 (jim_thompson5910):

just shorthand notation

OpenStudy (anonymous):

Thats cool, can i ask you other question?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

One more thing, so this is homework called webwork online Homework, I entered the answer it says log is too many inputs what does it mean

jimthompson5910 (jim_thompson5910):

show me with a screenshot please

OpenStudy (anonymous):

screenshot?

jimthompson5910 (jim_thompson5910):

take a picture of the screen using the printscreen key so I can see the error message

OpenStudy (anonymous):

sure,

jimthompson5910 (jim_thompson5910):

ok so the log isn't in the exponent at all

jimthompson5910 (jim_thompson5910):

you would use the product rule in this case

jimthompson5910 (jim_thompson5910):

y = 4^x*log(7,x) y' = 4^x*ln(4)*log(7,x) + 4^x*1/x*1/ln(7) y ' = (4^x*ln(4)*ln(x))/ln(7) + (4^x)/(x*ln(7))

jimthompson5910 (jim_thompson5910):

you can optionally factor out the GCF to get y ' = ( (4^x)/ln(7) ) * (ln(4)*ln(x) + 1/x )

OpenStudy (anonymous):

YES! you are RIGHT!. I have been asking this for about 4 hrs and people helped me but it didnt work. THANK YOU. can i ask 5 more questions ? its homework due tomorrow.

OpenStudy (anonymous):

i will open up the new post so you can have medal?

jimthompson5910 (jim_thompson5910):

the system didn't like the log(7,x) since it probably wanted ln(x)/ln(7) or log(x)/log(7)

jimthompson5910 (jim_thompson5910):

but I think log(7,x) is much more compact

OpenStudy (anonymous):

Yes, Thank you so much! Do you have time for other questions?

jimthompson5910 (jim_thompson5910):

just ask them here, it's fine

jimthompson5910 (jim_thompson5910):

yeah if they're not too long

jimthompson5910 (jim_thompson5910):

5 shouldn't be too bad

OpenStudy (anonymous):

I will delete this. post new one